OpenStudy (anonymous):

WILL FAN AND MEDAL Which choice is equivalent to the expression below? 2^6.38 A 2^(6+3/10+8/10) B 2^6*2^3/10*2^8/100 C 2^6*2^38/10 D 2^6+2^3/10+2^8/100

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

6.38=6+0.3+0.08 ths means that 2*6.38=2*(6+0.3+0.08)

OpenStudy (anonymous):

I meant 2^6.38=2^(6+0.3+0.08)

OpenStudy (anonymous):

2^(6+3/10+8/100)=2^6*2^(3/10)*2^(8/100)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@Dianie